OpenStudy (anonymous):

if I have to plot sqrt 5 on a graph how do i know where to plot it?

OpenStudy (amistre64):

make a graph of sqrt(x) and when x=5, go up to the y value.... calculators are useful for figuring out the sqrt(5)....it can be done by handm but might take awhile

OpenStudy (anonymous):

\(4 < 5 < 9 \implies \sqrt{4} < \sqrt{5} < \sqrt{9} \implies 2 < \sqrt{5} < 3\)

OpenStudy (anonymous):

So that should give you some ballpark.

OpenStudy (amistre64):

sqrt(4)<sqrt(5) < sqrt(9)...yeah like polpak did ;)

OpenStudy (anonymous):

And since 5 is a lot closer to 4 than it is to 9, the sqrt(5) will be closer to 2 than to 3.

OpenStudy (anonymous):

by hand: draw a right angle triangle with side 1 and 2, so hypotenuse is exactly sqrt(5), the use ur ruler :D

OpenStudy (amistre64):

ttt is still in kindergarten ;)

OpenStudy (anonymous):

oh yeah, miss teacher is too sexy that I dont want to leave kindergarten =))

OpenStudy (anonymous):

k i have . thanks!